# Do Language Models Understand the Cognitive Tasks Given to Them? Investigations with the N-Back Paradigm

### v2

@@ Abstract
Cognitive tasks originally developed for humans are now increasingly used to study language models. While applying these tasks is often straightforward, interpreting their results can be challenging. In particular, when a model underperforms, it is often unclear whether this results from a limitation in the cognitive ability being tested or a failure to understand the task itself. A recent study argues that GPT 3.5’s declining performance on 2-back and 3-back tasks reflects a working memory capacity limit similar to humans (Gong et al., 2024). By analyzing a range of open-source language models of varying performance levels on these tasks, we show that the poor performance instead reflects a limitation in task comprehension and task set maintenance. In addition, we challenge the best-performing model with progressively harder versions of the task (up to 10-back) and experiment with alternative prompting strategies, before analyzing model attentions. Our larger aim is to contribute to the ongoing conversation around refining methodologies for the cognitive evaluation of language models.
@@ 1 Introduction
A recent study applies the n-back task (Figure 1) to GPT 3.5 and concludes from the model’s poor 2-back and 3-back performance that it has a working memory capacity limit of approximately 3, apparently similar to humans (Gong et al., 2024). This interpretation raises two concerns. First, while working memory capacity limits are well established in human cognition, we cannot assume these same constraints exist or can be meaningfully measured in language models. Second, these results may simply reflect the model’s failure to understand the task requirements rather than any inherent memory limitation.
In this paper, we show that low-performing language models, even when provided with detailed n-back task instructions and demonstrations, commit errors that are consistent with a different m-back task. Notice that, if a human subject committed such systematic errors, we would conclude that they had misunderstood the task. In comparison, high-performing models consistently execute the correct task, while intermediate models tend to start with the correct task but drift toward a different one as errors accumulate.
The remainder of this paper is organized as follows. Section 2 covers relevant background. Section 3 introduces the dataset, prompting approach, and evaluation metrics. In Section 4.1, we evaluate a range of language models on the n-back task, focusing on the stimuli retrieved for comparison with the current ones. In Section 4.2, we show that, while successful models correctly map n-back instructions to n-back consistent retrievals (i.e., assign the highest log probabilities to n-back consistent retrievals when given n-back instructions), weaker models would map 2-back and 3-back instructions to 1-back consistent retrievals. We also analyze the effect of task instructions and demonstrations on these mappings. In Section 4.3, we find that, while the intermediate models are initially able to follow 2-back and 3-back instructions, as more stimuli come in, they tend to drift toward 1-back and 2-back consistent retrievals, respectively. In Section 4.4, we challenge the best model to perform 1-back through 10-back tasks and notice a signature of task comprehension: the model assigns the highest log probabilities to correct n-back continuations while assigning progressively lower log probabilities to m-back continuations as m deviates from n. In Sections 4.5 to 4.7, we discuss additional experiments with alternative prompting strategies for comparison. In Section 4.8, we identify an attention pattern whose prevalence predicts 2-back task performance.
@@ 2 Background and Related Work
There has been a growing body of work running cognitive tasks on pre-trained language models. These efforts often aim to identify whether the models exhibit cognitive constructs or capabilities that are present in humans. Subjects of study include theory of mind (Strachan et al., 2024; Gandhi et al., 2024), analogical reasoning (Hu et al., 2023; Webb et al., 2023), cognitive biases (Binz and Schulz, 2023; Lampinen et al., 2024), and working memory capacity (Gong et al., 2024), among many others.
@@
Such evaluations are susceptible to both overclaiming and underclaiming. On the one hand, false positives can result from training data contamination (Sainz et al., 2023), potentially compromising the validity of vignette-based assessments where models may produce memorized responses. On the other hand, underestimation of model capabilities can happen when we erroneously assume task comprehension, especially for smaller models (Hu and Frank, 2024). Prior studies have also investigated how well language models adhere to prompt instructions (Webson and Pavlick, 2022) and how their performance compares to that of humans (Webson et al., 2023). In light of other methodological challenges in the cognitive evaluation of language models, such as prompt sensitivity and cultural biases, Ivanova 2023 outlines recommendations for best practices.
@@
Virtually any task, from routine text comprehension to complex problem solving, involves the creation of intermediate or partial results. Successful task completion requires that these results be maintained in a way that facilitates later access. In humans, this mechanism is known as working memory, one of the most studied constructs in psychology for over half a century (Miyake and Shah, 1999). This concept can be extended to transformer-based language models designed to process interdependent, serial information. In fact, the transformer architecture, particularly its attention mechanism where key-query matching drives retrieval (Vaswani et al., 2017), bears striking resemblance to cue-based parsing and retrieval models proposed in psycholinguistics (Lewis et al., 2006), making it a promising candidate for modeling human sentence processing.
@@
One of the most salient and mysterious aspects of human working memory is its severely constrained capacity (Miller, 1956; Cowan, 2012). This capacity limit is often measured using the n-back task (Kirchner, 1958). To the best of our knowledge, Gong et al. 2024 is the first to apply the n-back task to a language model, specifically the GPT 3.5 Turbo variant of ChatGPT. They experiment with different prompting strategies, including those incorporating feedback and reasoning. As n increases from 1 to 3, they observe a sharp decline in model performance and conclude that ChatGPT has a working memory capacity limit of approximately 3.
@@
Figure 2: Retrieval accuracies on 1-back, 2-back, and 3-back tasks, averaged across 50 trials, for all models tested, organized by model family.
@@ 3.1 Data and Prompts
We use the same dataset as Gong et al. 2024. For each n-back task, there are 50 trials in total. Each trial consists of a sequence of 24 letters. In exactly 8 random positions within each sequence, the letters are the same as those appearing n position(s) earlier.
After each letter prompt, the models are instructed to answer “[ current letter ] and [ letter n back ] are [ different / identical ]”. This is designed to facilitate chain-of-thought reasoning (Wei et al., 2022) and to make explicit the specific letter retrieved by the model for comparison with the current one.
To teach the models the correct answer format and maximize their chances of correctly inferring the tasks, each trial begins with a demonstration, which includes a sequence of 24 letters and the correct responses. The “without demo” trials in Section 4.2 are the only exception. Following the demonstration, a new sequence of 24 letters is presented, one at a time, and the models are prompted to respond after each letter. An example 1-back trial is shown below; actual model responses are underlined.
$\texttt{SYS: }&\texttt{[TASK INSTRUCTIONS]}\\[-2.41109pt] \texttt{USR: }&\texttt{k}\\[-2.41109pt] \texttt{LLM: }&\texttt{k and none are different.}\\[-2.41109pt] \texttt{USR: }&\texttt{k}\\[-2.41109pt] \texttt{LLM: }&\texttt{k and k are identical.}\\[-2.41109pt] \texttt{USR: }&\texttt{a}\\[-2.41109pt] \texttt{LLM: }&\texttt{a and k are different.}\\[-2.41109pt] &\vdots\end{aligned}\right\}\texttt{DEMO}\\[3.44444pt] &\left.\begin{aligned} \texttt{SYS: }&\texttt{[TASK INSTRUCTIONS]}\\[-2.41109pt] \texttt{USR: }&\texttt{e}\\[-2.41109pt] \texttt{LLM: }&\underline{\texttt{e and none are different.}}\\[-2.41109pt] \texttt{USR: }&\texttt{f}\\[-2.41109pt] \texttt{LLM: }&\underline{\texttt{f and e are different.}}\\[-2.41109pt] \texttt{USR: }&\texttt{f}\\[-2.41109pt] \texttt{LLM: }&\underline{\texttt{f and f are identical.}}\\ &\vdots\end{aligned}\right\}\texttt{TEST}\end{aligned}$
@@ 3.3 Metrics
The n-back task requires continuously matching the current letter and the letter from n steps back to determine the correct label. However, compared to binary labels, the retrieved letters can tell us more about the models’ understanding of the task. And since the correct label is almost always assigned given the correct retrieval, our analyses focus on the retrieval accuracies and the log probabilities of the retrieved letters.
But how can we be sure that a model has inferred the right task from the instructions? One hypothesis is that, despite being prompted to do the n-back task, the model might be following m-back instructions instead. To investigate this, we adopt counterfactual measures by providing n-back instructions and evaluating the accuracies and log probabilities of retrievals consistent with the m-back task. We also apply variants of these measures, which we detail in later sections.
4 Experimental Results
4.1 Task Performance
We begin by comparing retrieval accuracies across models for all three tasks (Figure 2) and categorizing them into three performance tiers (Table 1):
  • T3 models achieve nearly perfect retrieval accuracies on 1-back trials, but their performance drops to around 20% or lower on 2-back and 3-back trials.
  • T2 models achieve nearly perfect retrieval accuracies on 1-back trials and around 50% and 40% on 2-back and 3-back trials, respectively.
  • T1 models achieve 100% retrieval accuracies on 1-back trials and over 80% on 2-back and 3-back trials.
@@ 4.1 Task Performance
Table 1: Retrieval accuracies on 3-back, 2-back, and 1-back tasks, averaged across 50 trials, for all models tested, organized by performance tier.
@@ 4.2 Task Comprehension
To better understand the source of these performance disparities, we ask: are less successful models able to infer the task from the provided instructions and demonstrations? Moreover, are high-performing models relying more heavily on task cues from the instructions or demonstrations? To address these questions, we:
  1. 1.
    Provide n-back task instructions and/or demonstrations.
  2. 2.
    Present three continuations, each consistent with a different m-back task.
  3. 3.
    Measure the average log probabilities of letters at retrieval positions for each trial.
@@
Under 1-back instructions, $\mathsf{P}_{1,1}>\mathsf{P}_{1,2}>\mathsf{P}_{1,3}$ across all models. The same is true when no task demonstrations are provided, with no significant difference between $\mathsf{P}_{1,m}$ and $\mathsf{P}^{-}_{1,m}$ for $m=1,2,3,$ as shown in Figure 3. Overall, this is unsurprising, given the near-perfect performance of all models on 1-back trials.
@@
  • T3.
    Under 2-back instructions, including with demonstrations, the model assigns 1-back continuations to be the most plausible, with both $\mathsf{P}^{-}_{2,1}>\mathsf{P}^{-}_{2,2}>\mathsf{P}^{-}_{2,3}$ and $\mathsf{P}_{2,1}>\mathsf{P}_{2,2}>\mathsf{P}_{2,3}$. The task demonstrations do bring $\mathsf{P}_{2,2}$ and $\mathsf{P}_{2,3}$ closer to $\mathsf{P}_{2,1}$, although this is not enough to offset the strong 1-back priors.
  • T2.
    Under 2-back instructions only, the ordering of $\mathsf{P}^{-}_{2,m}$ remains the same, albeit with $\mathsf{P}^{-}_{2,2}$ and $\mathsf{P}^{-}_{2,3}$ noticeably closer to $\mathsf{P}^{-}_{2,1}$ than for the T3 model. However, this time, when provided with the additional task demonstrations, the model finally assigns 2-back continuations to be the most likely, with $\mathsf{P}_{2,2}>\mathsf{P}_{2,1}>\mathsf{P}_{2,3}$.
  • T1.
    Somewhat surprisingly, we notice that $\mathsf{P}^{-}_{2,2}>\mathsf{P}^{-}_{2,1}>\mathsf{P}^{-}_{2,3}$, showing that the model is able to infer the task from the instructions alone. However, the demonstrations do help further consolidate the mapping.
@@
Thus, through analyzing models from different performance tiers, we clearly identify three distinct levels of task comprehension capabilities:
  • T3.
    The model cannot map 2-back and 3-back instructions to the correct responses, given either the instructions or demonstrations.
  • T2.
    The model cannot map 2-back and 3-back instructions to the correct responses, given the instructions, but can do so if demonstrations are also provided.
  • T1.
    The model can map 2-back and 3-back instructions to the correct responses based on the instructions alone, although this mapping is augmented by further demonstrations.
Figure 6: $\mathsf{A}_{2,\cdot}(i,m)$ for $m=1,2$ and $3\leq i\leq 24$. From top to bottom are results for Qwen 1.5 14b Chat (T3), Gemma 2 27b Instruct (T2), and Llama 3.1 70b Instruct (T1).
Figure 7: $\mathsf{A}_{3,\cdot}(i,m)$ for $m=1,2,3$ and $4\leq i\leq 24$. From top to bottom are results for Qwen 1.5 14b Chat (T3), Gemma 2 27b Instruct (T2), and Llama 3.1 70b Instruct (T1).
@@ 4.3 Task Set Maintenance
Each n-back trial presents a sequence of 24 letters. Successful task completion requires consistent adherence to the task instructions as more stimuli are presented. Here, we investigate whether language models show a progressive decline in their ability to produce n-back consistent responses over time. Previously, performance metrics were averaged across time steps for each trial; now, we average across trials for each time step.
Specifically, at each time step $i$ in the n-back task, we measure the average accuracy of m-back consistent retrievals for each $m\leq n$, given the model’s own responses up to time step $i-1$. Denote this as $\mathsf{A}_{n,\cdot}(i,m)$.
@@
For 1-back trials, $\mathsf{A}_{1,\cdot}(i,1)$ stays close to 1 for each model as $i$ increases (not shown), which is unsurprising.
@@
  • T3.
    Throughout the task, $\mathsf{A}_{2,\cdot}(i,1)$ stays close to $1$ and $\mathsf{A}_{2,\cdot}(i,2)$ stays close to 00. This is consistent with findings from Section 4.2.
  • T2.
    At first, the model tends to perform the right task. Over time, however, $\mathsf{A}_{2,\cdot}(i,2)$ decreases while $\mathsf{A}_{2,\cdot}(i,1)$ increases, with the latter eventually overtaking the former halfway through the task. In other words, the model is initially able to follow 2-back instructions, but, over time, the gradual accumulation of 1-back consistent errors ultimately shifts the model’s behavior away from the intended task.
  • T1.
    Throughout the task, $\mathsf{A}_{2,\cdot}(i,2)$ stays close to $1$ and $\mathsf{A}_{2,\cdot}(i,1)$ stays close to 00, which is the exact opposite of T3.
@@
  • T3.
    Throughout the task, $\mathsf{A}_{3,\cdot}(i,1)$ stays close to $1$ while both $\mathsf{A}_{3,\cdot}(i,2)$ and $\mathsf{A}_{3,\cdot}(i,3)$ stay close to 0, consistent with Section 4.2.
  • T2.
    After a transient initial lead, $\mathsf{A}_{3,\cdot}(i,3)$ is quickly surpassed by $\mathsf{A}_{3,\cdot}(i,2)$, suggesting yet greater difficulty with task set maintenance.
  • T1.
    Throughout the task, $\mathsf{A}_{3,\cdot}(i,3)$ remains close to 1, though it shows a gradual decline over time. Meanwhile, $\mathsf{A}_{3,\cdot}(i,1)$ and $\mathsf{A}_{3,\cdot}(i,2)$ remain relatively close to each other, with $\mathsf{A}_{3,\cdot}(i,2)$ slowly rising to be slightly greater than $\mathsf{A}_{3,\cdot}(i,1)$.
Effect of error accumulation.
Despite 2-back instructions and demonstrations, the T2 model gradually drifts toward 1-back consistent responses over time, suggesting that the accumulation of 1-back consistent errors may have significantly biased subsequent responses. To test this hypothesis, we manipulate the model’s response history by providing m-back consistent responses for i steps following n-back instructions and demonstrations, then compute the average m-back accuracy for time steps $i+1$ through 24, which we denote as $\mathsf{A}_{n,m}(i+1:24,m)$. Figure 8 shows that, as 1-back errors accumulate, 1-back responses are increasingly favored by the T2 model for subsequent steps, despite 2/3-back instructions and demonstrations. In comparison, both $\mathsf{A}_{2,2}(i+1:24,2)$ and $\mathsf{A}_{3,3}(i+1:24,3)$ remain relatively low, showing that correct responses do not bias subsequent answers to the same degree.
Figure 8: Top: $\mathsf{A}_{2,m}(i+1:24,m)$ for $m=1,2$ and $3\leq i\leq 23$, using Gemma 2 27b Instruct (T2). Bottom: $\mathsf{A}_{3,m}(i+1:24,m)$ for $m=1,2,3$ and $4\leq i\leq 23$, using the same model.

### v3

# Instructions for *ACL Proceedings

@@ Abstract
Cognitive tasks originally developed for humans are now increasingly used to study language models. While applying these tasks is often straightforward, interpreting their results can be challenging. In particular, when a model underperforms, it is often unclear whether this results from a limitation in the cognitive ability being tested or a failure to understand the task itself. A recent study argues that GPT 3.5’s declining performance on 2-back and 3-back tasks reflects a working memory capacity limit similar to humans (Gong et al., 2024). By analyzing a range of open-source language models of varying performance levels on these tasks, we show that the poor performance is due at least in part to a limitation in task comprehension and task set maintenance. We challenge the best-performing model with progressively harder versions of the task (up to 10-back) and experiment with alternative prompting strategies, before analyzing model attentions. Our larger aim is to contribute to the ongoing conversation around refining methodologies for the cognitive evaluation of language models. [^1] Code available at https://github.com/hxiaoyang/lm-nback.
Do Language Models Understand the Cognitive Tasks Given to Them?
Investigations with the N-Back Paradigm
Xiaoyang Hu Brown University xiaoyang_hu@brown.edu                        Richard L. Lewis University of Michigan rickl@umich.edu
@@ 1 Introduction
A recent study applies the $n$-back task (Figure 1) to GPT 3.5 and concludes from the model’s poor 2-back and 3-back performance that it has a working memory capacity limit (WMCL) of approximately 3, apparently similar to humans (Gong et al., 2024). This interpretation raises two concerns. First, while WMCL is well established in human cognition, we cannot assume these same constraints exist or can be meaningfully measured in language models. Second, these results may reflect the model’s failure to understand the task requirements rather than any inherent memory limitation.
In this paper, we show that low-performing language models, even when provided with detailed $n$-back task instructions and demonstrations, commit errors that are consistent with a different $m$-back task ($m\neq n$). Notice that, if a human subject committed such systematic errors, we would conclude that they had misunderstood the task. In comparison, intermediate models, including GPT 3.5, tend to start with the correct task but drift toward a different one as errors accumulate, resulting in poor average 2-back and 3-back performance, consistent with Gong et al. 2024. High-performing models, on the other hand, consistently execute the correct task, even for larger $n$’s, achieving task accuracies of 90.08%, 90.08%, and 84.75% for $n=8,9,10$.
The remainder of this paper is organized as follows. Section 2 covers relevant background and related work. Section 3 introduces the dataset, models, prompting approach, and evaluation metrics. Section 4.1 benchmarks each model on 1-back, 2-back, and 3-back tasks, identifying three distinct performance tiers. Section 4.2 investigates whether these performance disparities are explained by differences in task comprehension. Section 4.3 examines the models’ ability to consistently apply the correct task set throughout each trial (task set maintenance). In Section 4.4, we challenge the best model to perform 1-back through 10-back tasks and notice a signature of task comprehension. In Sections 4.5 and 4.6, we discuss additional experiments with alternative prompting strategies for comparison. In Section 4.7, we identify an attention pattern whose prevalence predicts 2-back task performance.
@@ 2 Background and Related Work
There has been a growing body of work that evaluates pre-trained language models using cognitive tasks originally developed for humans. These efforts often aim to identify whether the models exhibit cognitive constructs or capabilities that are present in humans. Subjects of study include theory of mind (Strachan et al., 2024; Gandhi et al., 2024), analogical reasoning (Hu et al., 2023; Webb et al., 2023), cognitive biases (Binz and Schulz, 2023; Lampinen et al., 2024), and WMCL (Gong et al., 2024), among many others. Such evaluations are susceptible to both overclaiming and underclaiming. On the one hand, false positives can result from training data contamination (Sainz et al., 2023), potentially compromising the validity of vignette-based assessments where models may produce memorized responses. On the other hand, underestimation of model capabilities can happen when we erroneously assume task comprehension, especially for smaller models (Hu and Frank, 2024). Prior studies have also investigated how well language models adhere to prompt instructions, especially compared to humans (Webson and Pavlick, 2022; Webson et al., 2023). In light of other methodological challenges in the cognitive evaluation of language models, such as prompt sensitivity and cultural biases, Ivanova 2023 outlines recommendations for best practices.
@@
Virtually any task, from routine text comprehension to complex problem solving, involves the creation of intermediate or partial results. Successful task completion requires that these results be maintained in a way that facilitates later access. In humans, this mechanism is known as working memory, one of the most studied constructs in psychology for over half a century (Miyake and Shah, 1999). This concept can be extended to transformer-based language models designed to process interdependent, serial information. In fact, the transformer architecture, particularly its attention mechanism where key-query matching drives retrieval (Vaswani et al., 2017), bears striking resemblance to cue-based parsing and retrieval models proposed in psycholinguistics (Lewis et al., 2006), making it a promising candidate for modeling human sentence processing. One of the most salient and mysterious aspects of human working memory is its severely constrained capacity (Miller, 1956; Cowan, 2012). One prominent task used to measure working memory capacity is the $n$-back task (Kirchner, 1958).
@@
To the best of our knowledge, Gong et al. 2024 are the first to apply the $n$-back task to a language model, specifically the GPT 3.5 Turbo variant of ChatGPT. They experiment with different prompting strategies, including those incorporating feedback and reasoning. As $n$ increases from 1 to 3, they observe a sharp decline in model performance and conclude that the model has a WMCL of approximately 3. Zhang et al. 2024 also examine working memory in language models using a task described as $n$-back, but with all stimuli presented simultaneously. This departs from the standard paradigm and imposes different working memory demands. Moreover, while they acknowledge that the poor performance in smaller models may stem from limited understanding of the “intent of the input”, they do not control for task comprehension as a confounding variable.
Multi-hop question answering is another working memory task paradigm, in which two or more reasoning steps must be performed sequentially to resolve complex queries. This task is interestingly different from $n$-back in that it places implicit demands on working memory in a single forward pass. For instance, answering “The spouse of the performer of Imagine was…” requires first identifying John Lennon as the performer of the song and then determining that Yoko Ono was his spouse. Biran et al. 2024 find that, after early model layers resolve the initial step, later layers often prove deficient in completing the second.
@@
Figure 2: Average retrieval accuracies on 1-, 2-, and 3-back tasks, grouped by model family.
@@ 3.1 Data and Prompts
We use the dataset from Gong et al. 2024 (MIT License). For each $n$-back task, there are 50 trials in total. Each trial consists of a sequence of 24 letters. In exactly 8 random positions within each sequence, the letters are the same as those appearing $n$ step(s) earlier. After each letter prompt, the models are instructed to answer “{current letter} and {letter $n$ back} are {different / identical}” This is designed to facilitate chain-of-thought reasoning (Wei et al., 2022) and to make explicit the specific letter retrieved by the model for comparison with the current one.
$\texttt{SYS: }&\texttt{[TASK INSTRUCTIONS]}\\[-3.01385pt] \texttt{USR: }&\texttt{k}\\[-3.01385pt] \texttt{LLM: }&\texttt{k and none are different.}\\[-3.01385pt] \texttt{USR: }&\texttt{k}\\[-3.01385pt] \texttt{LLM: }&\texttt{k and k are identical.}\\[-3.01385pt] \texttt{USR: }&\texttt{a}\\[-3.01385pt] \texttt{LLM: }&\texttt{a and k are different.}\\[-3.01385pt] &\vdots\end{aligned}\right\}\texttt{DEMO}\\[4.30554pt] &\left.\begin{aligned} \texttt{SYS: }&\texttt{[TASK INSTRUCTIONS]}\\[-3.01385pt] \texttt{USR: }&\texttt{e}\\[-3.01385pt] \texttt{LLM: }&\underline{\texttt{e and none are different.}}\\[-3.01385pt] \texttt{USR: }&\texttt{f}\\[-3.01385pt] \texttt{LLM: }&\underline{\texttt{f and e are different.}}\\[-3.01385pt] \texttt{USR: }&\texttt{f}\\[-3.01385pt] \texttt{LLM: }&\underline{\texttt{f and f are identical.}}\\ &\vdots\end{aligned}\right\}\texttt{TEST}\end{aligned}$
To teach the models the correct answer format and maximize their chances of correctly inferring the tasks, each trial begins with a demonstration, which includes a sequence of 24 letters and the correct responses. The “without demo” trials in Section 4.2 are the only exception. Following the demonstration, a new sequence of 24 letters is presented, one at a time, and the models are prompted to respond after each letter. An example 1-back trial is shown above; actual model responses are underlined.
@@ 3.3 Metrics
The $n$-back task requires continuously matching the current letter and the letter from $n$ steps back to determine the correct label. Compared to binary labels, the retrieved letters offer better insight into the models’ understanding of the task. And since the correct label is almost always assigned given the correct retrieval, our analyses focus on the retrieval accuracies and the log probabilities of the retrieved letters. One possibility for low performance is that, despite being prompted to do the $n$-back task, a model might be following $m$-back instructions instead. To investigate this, we adopt counterfactual measures by providing $n$-back instructions and evaluating the accuracies and log probabilities of retrievals consistent with the $m$-back task. We also apply variants of these measures, which we detail in later sections.
@@
Table 1: Average retrieval accuracies on 1-, 2-, and 3-back tasks, organized by performance tier.
4 Experimental Results
4.1 Task Performance
We begin by comparing retrieval accuracies across models for all three tasks (Figure 2) and categorizing them into three performance tiers (Table 1).
T3: $\leq 20\%$ on 2- and 3-back.
T2: $\sim 50\%$ on 2-back; $\sim 40\%$ on 3-back.
T1: $>80\%$ on 2- and 3-back.
@@ 4.2 Task Comprehension
To better understand the source of these performance disparities, we ask: Can models infer the task from the instructions and demonstrations—and if so, which serves as a more effective cue? To address these questions, we 1) provide $n$-back instructions with and without demonstrations, 2) present three continuations, each consistent with a different $m$-back task, and 3) measure the average retrieval log probabilities for each trial.
@@
Under 1-back instructions, $\mathsf{P}_{1,1}>\mathsf{P}_{1,2}>\mathsf{P}_{1,3}$ across all models. The same is true when no task demonstrations are provided, with no significant difference between $\mathsf{P}_{1,m}$ and $\mathsf{P}^{-}_{1,m}$ for $m=1,2,3,$ as shown in Figure 3. Overall, this is unsurprising, given the near-perfect performances of all models on 1-back trials.
@@
T3: Under 2-back instructions, including with demonstrations, 1-back continuations are assigned to be the most plausible, with both $\mathsf{P}^{-}_{2,1}>\mathsf{P}^{-}_{2,2}>\mathsf{P}^{-}_{2,3}$ and $\mathsf{P}_{2,1}>\mathsf{P}_{2,2}>\mathsf{P}_{2,3}$. The task demonstrations do bring $\mathsf{P}_{2,2}$ and $\mathsf{P}_{2,3}$ closer to $\mathsf{P}_{2,1}$, although this is not enough to offset the strong 1-back priors.
T2: Under 2-back instructions only, the ordering of $\mathsf{P}^{-}_{2,m}$ remains the same, albeit with $\mathsf{P}^{-}_{2,2}$ and $\mathsf{P}^{-}_{2,3}$ noticeably closer to $\mathsf{P}^{-}_{2,1}$ than for T3. However, with additional task demonstrations, 2-back continuations are assigned to be the most likely, with $\mathsf{P}_{2,2}>\mathsf{P}_{2,1}>\mathsf{P}_{2,3}$.
T1: Somewhat surprisingly, we notice that $\mathsf{P}^{-}_{2,2}>\mathsf{P}^{-}_{2,1}>\mathsf{P}^{-}_{2,3}$, showing that the model is able to infer the task from the instructions alone. However, the demonstrations do help further consolidate the mapping.
@@
Summary.
Through analyzing models from different performance tiers, we identify three distinct levels of task comprehension capabilities. The T3 model fails to map 2-back and 3-back instructions to the correct responses, given either the instructions or demonstrations, suggesting it completely misunderstands the task; the T2 model fails to map 2-back and 3-back instructions to the correct responses, given the instructions, but can do so if demonstrations are also provided; the T1 model can map 2-back and 3-back instructions to the correct responses based on the instructions alone, suggesting a robust understanding of the tasks. The T3 model’s failure to understand the task is corroborated by analyses in Section 4.6. Even when provided with short demo sequences and immediate corrective feedback, it still fails to get 2 consecutive correct responses. This suggests that its poor performance stems from an inability to understand the task, rather than any memory limitation.
@@ 4.3 Task Set Maintenance
Each $n$-back trial consists of a sequence of 24 letters. Successful task completion requires consistent adherence to the task instructions as more stimuli are presented. Here, we investigate whether language models show a progressive decline in their ability to produce $n$-back consistent responses over time. Previously, performance metrics were averaged across time steps for each trial. Now, we average across trials for each time step. At each time step $i$ in the $n$-back task, we measure the average accuracy of $m$-back consistent retrievals for each $m\leq n$, given the model’s own responses up to time step $i-1$. Denote this as $\mathsf{A}_{n,\cdot}(m,i)$.
Figure 6: $\mathsf{A}_{2,\cdot}(m,i)$ for $m=1,2$ and $3\leq i\leq 24$. From top to bottom are results for Qwen 1.5 14b Chat (T3), Gemma 2 27b Instruct (T2), and Llama 3.1 70b Instruct (T1).
Figure 7: $\mathsf{A}_{3,\cdot}(m,i)$ for $m=1,2,3$ and $4\leq i\leq 24$. From top to bottom are results for Qwen 1.5 14b Chat (T3), Gemma 2 27b Instruct (T2), and Llama 3.1 70b Instruct (T1).
@@
Unsurprisingly, $\mathsf{A}_{1,\cdot}(1,i)$ stays close to 1 for each model as $i$ increases (not shown).
@@
T3: Throughout the task, $\mathsf{A}_{2,\cdot}(1,i)$ and $\mathsf{A}_{2,\cdot}(2,i)$ stay close to $1$ and 00, respectively, consistent with findings from Section 4.2.
T2: $\mathsf{A}_{2,\cdot}(2,i)$ crosses below $\mathsf{A}_{2,\cdot}(1,i)$ halfway through the task, suggesting a gradual shift from 2-back to 1-back behavior.
T1: Throughout the task, $\mathsf{A}_{2,\cdot}(2,i)$ and $\mathsf{A}_{2,\cdot}(1,i)$ stay close to $1$ and 00, respectively, contrary to T3.
@@
T3: Throughout the task, $\mathsf{A}_{3,\cdot}(1,i)$ stays close to $1$ while both $\mathsf{A}_{3,\cdot}(2,i)$ and $\mathsf{A}_{3,\cdot}(3,i)$ stay close to 0, consistent with Section 4.2.
T2: After a transient initial lead, $\mathsf{A}_{3,\cdot}(3,i)$ is quickly surpassed by $\mathsf{A}_{3,\cdot}(2,i)$, suggesting yet greater difficulty with task set maintenance.
T1: Throughout the task, $\mathsf{A}_{3,\cdot}(3,i)$ remains close to 1, though it shows a gradual decline over time. Meanwhile, $\mathsf{A}_{3,\cdot}(1,i)$ and $\mathsf{A}_{3,\cdot}(2,i)$ remain low.
Figure 8: Top: $\mathsf{A}_{2,m}(m,i+1:24)$ for $m=1,2$ and $3\leq i\leq 23$, using Gemma 2 27b Instruct (T2). Bottom: $\mathsf{A}_{3,m}(m,i+1:24)$ for $m=1,2,3$ and $4\leq i\leq 23$, using the same model.
@@
Effect of error accumulation.
Despite 2-back instructions and demonstrations, the T2 model gradually drifts toward 1-back consistent responses over time, suggesting that the accumulation of 1-back consistent errors may have significantly biased subsequent responses. To test this hypothesis, we manipulate the model’s response history by providing $m$-back consistent responses for i steps following $n$-back instructions and demonstrations. We then compute the average $m$-back accuracy for time steps $i+1$ through 24, denoted as $\mathsf{A}_{n,m}(m,i+1:24)$. Figure 8 shows that, as 1-back errors accumulate, 1-back responses are increasingly favored by the T2 model for subsequent steps, despite 2- or 3-back instructions and demonstrations. In comparison, both $\mathsf{A}_{2,2}(2,i+1:24)$ and $\mathsf{A}_{3,3}(3,i+1:24)$ remain relatively low, showing that correct responses do not bias subsequent answers to the same degree.